# Category-free Out-of-Distribution Node Detection with Feature Resonance

### v1

@@ 3.3 Extension with Synthetic OOD Node Strategy
Although the resonance-based filtering score effectively separates OOD nodes, recent studies (Gong & Sun, 2024) suggest that training an OOD classifier with synthetic OOD nodes can improve OOD node detection. Therefore, we propose a novel framework that employs feature resonance scores to generate more realistic synthetic OOD nodes.
Specifically, we define the candidate OOD node set as $\mathcal{V}_{\mathrm{cand}}=\{\tilde{v}_{i}\in\mathcal{V}_{\mathrm{wild}}:\tau_{i}\leq T\}$, where $T=\mathrm{min}_{\mathrm{n}}(\tau)$ is the $n$-th smallest $\tau$ of wild nodes, selecting nodes with the smallest $n$ $\tau$ values. The features of these nodes form $\boldsymbol{X}_{\mathrm{cand}}$. Then, we compute a trainable metric based on the weighted mapping of node $v$’s representations across $K$ GNN layers: $E_{\theta}(v)=\mathbf{W}_{K}\big{(}\sum_{k}^{K}\beta_{k}\mathbf{h}_{v}^{(k)}\big{)}$, where $\beta_{k}\in\mathbb{R}$ is a learnable parameter, and $\mathbf{W}_{K}\in\mathbb{R}^{1\times d}$ transforms the node representations to the energy scalar. Then, we employ stochastic gradient Langevin dynamics (SGLD) (Welling & Teh, 2011) to generate synthetic OOD nodes $\mathcal{V}_{\mathrm{syn}}=\{\hat{v}_{1},\cdots,\hat{v}_{j}\}$ with random initial features $\boldsymbol{X}_{\mathrm{syn}}=\{\hat{\mathbf{x}}_{1},\cdots,\hat{\mathbf{x}}_{j}\}$ as follows:
$\begin{split}\hat{\mathbf{x}}_{j}^{(t+1)}=&\lambda\big{(}\hat{\mathbf{x}}_{j}^{(t)}-\frac{\alpha}{2}\nabla_{\hat{\mathbf{x}}_{j}^{(t)}}E_{\theta}\big{(}\hat{v}_{j}^{(t)}\big{)}+\epsilon\big{)}\\ &+(1-\lambda)\mathbb{E}_{\mathbf{x}\sim\boldsymbol{X}_{\mathrm{cand}}}(\mathbf{x}-\hat{\mathbf{x}}_{j}^{(t)})\end{split}$ (6)
where $\frac{\alpha}{2}$ is the step size and $\lambda$ is a trade-off hyperparameter. $\epsilon$ is the Gaussian noise sampled from multivariate Gaussian distribution $\mathcal{N}(0,\zeta)$. Unlike EnergyDef (Gong & Sun, 2024), we utilize the candidate OOD nodes $\mathcal{V}_{\mathrm{cand}}$ as examples to generate synthetic OOD nodes that better align with the actual OOD nodes. After obtaining the synthetic OOD nodes, we define the training set $\mathcal{V}_{\mathrm{train}}=\mathcal{V}_{\text{known}}\ \cup\ \mathcal{V}_{\mathrm{cand}}\ \cup\ \mathcal{V}_{\mathrm{syn}}$ with features $\boldsymbol{X}_{\mathrm{train}}$ and labels $\boldsymbol{Y}_{\mathrm{train}}$. The initially known ID nodes $\mathcal{V}_{\mathrm{known}}$ are assigned a label of $1$. In contrast, the candidate OOD nodes $\mathcal{V}_{\mathrm{cand}}$ and the generated synthetic OOD nodes $\mathcal{V}_{\mathrm{syn}}$ are assigned a label of 00. We use binary cross-entropy loss for training:
@@
Table 1: The statistics of the real-world OOD node detection datasets. $\times$ denotes no available multi-category labels. Notably, even on Squirrel and WikiCS, we do not use any true label as well.
Dataset Squirrel WikiCS YelpChi Amazon Reddit
# Nodes 5,201 11,701 45,954 11,944 10,984
# Features 2,089 300 32 25 64
Avg. Degree 41.7 36.9 175.2 800.2 15.3
OOD node (%) 20.0 29.5 14.5 9.5 3.3
# Category 5 10 $\times$ $\times$ $\times$
Table 2: Category-free OOD detection on real-world datasets. “OOM” indicates out-of-memory, “TLE” means time limit exceeded, and “-” denotes inapplicability. Detectors with $\clubsuit$ use only node attributes, while $\spadesuit$ share RSL’s GNN backbone. Entropy-based methods with $\lozenge$ use true multi-category labels, and $\blacklozenge$ rely on K-means pseudo labels. Top results: 1st, 2nd.
@@
3.4 Theoretical Analysis
Theorem 3.3.
$\begin{split}\text{ERR}_{\text{out}}^{t}\leq&\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\\ &+O(\sqrt{\frac{d}{\pi^{2}n}})+O(\sqrt{\frac{\mathrm{max}\{d,\Delta_{\xi}^{\eta^{2}}/\pi^{2}\}}{\pi^{2}(1-\pi)m}})\end{split}$ (8)
Practical implications of Therorem 3.3. The above theorem states that under mild assumptions, the error $ERR_{\mathrm{out}}$ is upper bounded. If the following two regulatory conditions hold: 1) the sizes of the labeled ID $n$ and wild data $m$ are sufficiently large; 2) the optimal ID risk $R_{in}^{*}$ is small, then the upper bound is mainly depended on $T$ and $t$. We further study the main error of $T$ and $t$ which we defined as $\delta(T,t)$.
Theorem 3.4.
$\begin{split}\delta(T,t)&=\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\\ &\leq\frac{\epsilon}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\end{split}$ (9)
Practical implications of Therorem 3.4. Theorem 3.4 states that when the learning rate $\alpha$ is sufficiently small, the primary error $\delta(T,t)$ can approach zero if the difference $\zeta$ between the two data distributions $\mathbb{P}_{\text{wild}}$ and $\mathbb{P}_{\text{in}}$ is greater than a certain small value. Meanwhile, Theorem 3.4 also shows that the primary error $\delta(T,t)$ is inversely proportional to the learning rate $\alpha$ and the number of epochs ($t$). As the $t$ increases, the primary error $\delta(T,t)$ also increases, while a smaller learning rate $\alpha$ leads to a minor primary error $\delta(T,t)$. However, during training, there exists $t$ at which the error reaches its minimum.
4 Experiment
4.1 Experimental Setup
Datasets.
We conduct extensive experiments to evaluate RSL on five real-world OOD node detection datasets: Two multi-category datasets, Squirrel (Rozemberczki et al., 2021) and WikiCS (Mernyei & Cangea, 2020), and three binary classification fraud detection datasets: YelpChi (Rayana & Akoglu, 2015), Amazon (McAuley & Leskovec, 2013), and Reddit (Kumar et al., 2019). The statistics of these datasets are summarized in Table 1. We provide detailed dataset description in the Appendix F.3.
Figure 4: Performance of detecting OOD nodes with different metrics. $\tau$ represents the resonance-based score, the “Overall Trajectory” represents the total cumulative length of the training trajectory $\hat{F}(\tilde{\mathbf{x}}_{i})=\sum_{t}\tau_{i}$, and the “Sliding Window” refers to the cumulative $\tau$ within a window of width 10: $\hat{F}_{10}(\tilde{\mathbf{x}}_{i})=\sum^{t}_{t-10}\tau_{i}$.
Baselines.
We assess the performance of RSL against a diverse range of baseline methods spanning five categories: 1) Traditional outlier detection methods, including local outlier factor (Breunig et al., 2000) with $k$-nearest neighbors (LOF-KNN) and MLP autoencoder (MLPAE). 2) Graph-based outlier detection models, including GCN autoencoder (Kipf & Welling, 2016b), GAAN (Chen et al., 2020), DOMINANT (Ding et al., 2019), ANOMALOUS (Peng et al., 2018), and SL-GAD (Zheng et al., 2021). 3) Transformation-based outlier detection approaches, such as GOAD (Bergman & Hoshen, 2020) and NeuTral AD (Qiu et al., 2021). 4) Entropy-based detection techniques, including GKDE (Zhao et al., 2020), OODGAT (Song & Wang, 2022), and GNNSafe (Wu et al., 2023). 5) Category-free detection methods, including EnergyDef (Gong & Sun, 2024) and SSD (Sehwag et al., 2021). Details of baselines and implementation are in Appendix F.4 and F.5, respectively.
(a) EnergyDef
(b) Resonance-based Filter Score $\tau$ (Ours)
(c) RSL (Ours)
Figure 5: The score distribution of ID nodes and OOD nodes on Amazon obtained using different methods.
Metrics.
Following prior research on OOD node detection, we evaluate the detection performance using three widely recognized, threshold-independent metrics: AUROC ($\uparrow$), AUPR ($\uparrow$) and FPR95($\downarrow$). We provide a detailed metric description in the Appendix F.2.
4.2 Main Results
Table 2 presents the main experimental results of various methods across five public datasets. The traditional methods like LOF-KNN and MLPAE perform poorly across most datasets, particularly with high false positive rates. Graph-based methods such as GCNAE, GAAN, and DOMINANT show some decent results but generally lag behind RSL and EnergyDef. Entropy-based methods like OODGAT and GNNSafe perform well on datasets with multi-category label information (e.g., WikiCS) but struggle on datasets without such labels, like YelpChi. Overall, these methods tend to be less robust compared to RSL. Specifically, RSL achieves significant improvements on most datasets. On average, it improves by 5.24% and 19.26% in AUROC and AUPR, respectively, and reduces FPR95 by 18.51%. Notably, when additional multi-category label information is available, GNNSafe outperforms EnergyDef on WikiCS. However, even in the absence of multi-class label information, our method still outperforms GNNSafe, achieving a remarkable improvement of 13.33% in AUROC, 22.07% in AUPR, and a 35.43% reduction in FPR95. These results powerfully demonstrate the effectiveness of our method.
Method Dataset Target Squirrel WikiCS
[…]
EnergyDef - 64.15 37.40 91.77 70.22 60.10 83.17
RSL w/o classifier A random vector 61.52 38.96 90.18 79.15 78.65 70.38
4.3 Ablation Study
How effective is resonance-based filter score $\tau$?
The experimental results in the row labeled “RSL w/o classifier” of Table 2 show that using the raw resonance-based score $\tau$ to filter OOD nodes is already more effective than the SOTA method on most datasets. On the FPR95 metric, the resonance-based score achieves an average reduction of 10.93% compared to current SOTA methods. Notably, even when compared to methods that leverage additional multi-category label information, our approach continues to demonstrate a clear performance advantage. For example, on the WikiCS dataset, the resonance-based score reduces the FPR95 metric by 7.69% compared to GNNSafe.
How effective are the synthetic OOD nodes combined with the feature resonance score?
The experimental results in the row labeled “RSL w/o $\mathcal{V}_{\mathrm{syn}}$” of Table 2 show that after removing the synthetic OOD nodes, the performance of the trained OOD classifier declined to varying degrees. This indicates that synthetic OOD nodes enhance the generalization ability of the OOD classifier, allowing it to detect more OOD nodes more accurately. It is worth noting that our synthetic OOD nodes, generated by leveraging real OOD nodes selected using $\tau$, better align with real-world OOD scenarios and, therefore, outperform EnergyDef.
4.4 Effectiveness of Feature Resonance Score in Selecting OOD Nodes
@@
We aim to evaluate the performance of RSL when integrated with methods other than the resonance-based score for selecting reliable OOD nodes. To ensure fairness, we used the same parameters and selected the same number of OOD nodes. From a metric learning perspective, we computed the cosine similarity, Euclidean distance, and Mahalanobis distance between unknown nodes and the prototypes of known ID nodes, with smaller values indicating a higher likelihood of being OOD nodes. We also applied EnergyDef for OOD node selection. The results, presented in Table 3, show that, under the same conditions, the OOD nodes selected using $\tau$ are more reliable than those selected by the other methods.
4.5 Effectiveness of Different Scoring Strategies Based on Feature Resonance
We evaluate the effectiveness of three score design strategies based on feature resonance: the resonance-based score $\tau$, the global trajectory norm, and the sliding window accumulation (width 10). As shown in Figure 4, $\tau$ outperforms the other two scores on most datasets. The sliding window approach performs better than the global trajectory norm, with further details on width in Appendix G.3. This indicates that finer-grained information improves OOD node detection, so we select $\tau$ as the primary score for filtering OOD nodes in our method.
4.6 Feature Resonance with Different Target Vectors
We explore micro-level feature resonance using different target vectors through experiments on Squirrel and WikiCS datasets with true $N$-category labels. Based on neural collapse theory (Papyan et al., 2020; Zhou et al., 2022), we set $N$ target vectors that form a simplex equiangular tight frame  [^3] The definition of the simplex equiangular tight frame is introduced in Appendix G.1., maximizing separation. As shown in Table 4, the ”True multi-label” row demonstrates the effectiveness of this approach. Interestingly, even when random labels are assigned (the ”Multiple random vectors” row) or when all ID representations align with a fixed vector (the ”A random vector” row), unknown ID nodes still show larger $\tau$ than unknown OOD nodes, as seen in Table 4. These results suggest that feature resonance is label-independent, stemming from intrinsic relationships between ID node representations, making it suitable for category-free OOD detection.
Table 5: Time cost (s).
Method Dataset Squirrel WikiCS YelpChi Amazon Reddit
EnergyDef 10.94 27.11 76.51 33.81 26.44
RSL w/o classifier 5.25 4.03 5.41 5.75 3.71
RSL 11.54 17.53 74.83 36.33 38.23
4.7 Time Efficiency
4.8 Score Distribution Visualization
We visualize the score distributions of ID and OOD nodes on the Amazon dataset obtained using different methods, as shown in Figure 5. When using the resonance-based score (Figure 5 (b)), the majority of unknown ID nodes show more significant representation changes compared to unknown OOD nodes. This separation of OOD nodes already exceeds EnergyDef (Figure 5 (a)). After training with synthetic OOD nodes (Figure 5 (c)), the separation between the energy scores of ID and OOD nodes still improves compared to EnergyDef, which demonstrates the effectiveness of RSL.
5 Related Works
General OOD Detection Methods.

### v2

@@ 2.3 Extension with Synthetic OOD Node Strategy
Although the resonance-based filtering score effectively separates OOD nodes, recent studies (Gong and Sun, 2024) suggest that training an OOD classifier with synthetic OOD nodes can improve OOD node detection. Therefore, we propose a novel framework that employs feature resonance scores to generate more realistic synthetic OOD nodes.
Specifically, we define the candidate OOD node set as $\mathcal{V}_{\mathrm{cand}}=\{\tilde{v}_{i}\in\mathcal{V}_{\mathrm{wild}}:\tau_{i}\leq T\}$, where $T=\mathrm{min}_{\mathrm{n}}(\tau)$ is the $n$-th smallest $\tau$ of wild nodes, selecting nodes with the smallest $n$ $\tau$ values. The features of these nodes form $\bm{X}_{\mathrm{cand}}$. Then, we compute a trainable metric based on the weighted mapping of node $v$’s representations across $K$ GNN layers: $E_{\theta}(v)=\mathbf{W}_{K}\big(\sum_{k}^{K}\beta_{k}\mathbf{h}_{v}^{(k)}\big)$, where $\beta_{k}\in\mathbb{R}$ is a learnable parameter, and $\mathbf{W}_{K}\in\mathbb{R}^{1\times d}$ transforms the node representations to the energy scalar. Then, we employ stochastic gradient Langevin dynamics (SGLD) (Welling and Teh, 2011) to generate synthetic OOD nodes $\mathcal{V}_{\mathrm{syn}}=\{\hat{v}_{1},\cdots,\hat{v}_{j}\}$ with random initial features $\bm{X}_{\mathrm{syn}}=\{\hat{\mathbf{x}}_{1},\cdots,\hat{\mathbf{x}}_{j}\}$ as follows:
$\begin{split}\hat{\mathbf{x}}_{j}^{(t+1)}=&\lambda\big(\hat{\mathbf{x}}_{j}^{(t)}-\frac{\alpha}{2}\nabla_{\hat{\mathbf{x}}_{j}^{(t)}}E_{\theta}\big(\hat{v}_{j}^{(t)}\big)+\epsilon\big)+(1-\lambda)\mathbb{E}_{\mathbf{x}\sim\bm{X}_{\mathrm{cand}}}(\mathbf{x}-\hat{\mathbf{x}}_{j}^{(t)})\end{split}$ (4)
where $\frac{\alpha}{2}$ is the step size and $\lambda$ is a trade-off hyperparameter. $\epsilon$ is the Gaussian noise sampled from multivariate Gaussian distribution $\mathcal{N}(0,\zeta)$. Unlike EnergyDef (Gong and Sun, 2024), we utilize the candidate OOD nodes $\mathcal{V}_{\mathrm{cand}}$ as examples to generate synthetic OOD nodes that better align with the actual OOD nodes. After obtaining the synthetic OOD nodes, we define the training set $\mathcal{V}_{\mathrm{train}}=\mathcal{V}_{\text{known}}\ \cup\ \mathcal{V}_{\mathrm{cand}}\ \cup\ \mathcal{V}_{\mathrm{syn}}$ with features $\bm{X}_{\mathrm{train}}$ and labels $\bm{Y}_{\mathrm{train}}$. The initially known ID nodes $\mathcal{V}_{\mathrm{known}}$ are assigned a label of $1$. In contrast, the candidate OOD nodes $\mathcal{V}_{\mathrm{cand}}$ and the generated synthetic OOD nodes $\mathcal{V}_{\mathrm{syn}}$ are assigned a label of 0. We use binary cross-entropy loss for training:
@@
Table 1: The statistics of the real-world OOD node detection datasets. $\times$ denotes no available multi-category labels. Notably, we do not use any true labels for all datasets.
Dataset Squirrel WikiCS Cora Citeseer Pubmed Chameleon YelpChi Amazon Reddit
# Nodes 5,201 11,701 2,708 3,327 19,717 2,277 45,954 11,944 10,984
# Features 2,089 300 1,433 3,703 500 2,325 32 25 64
Avg. Degree 41.7 36.9 7.8 5.5 9.0 31.7 175.2 800.2 15.3
OOD node (%) 20.0 29.5 66.7 45.7 20.8 40.2 14.5 9.5 3.3
# Category 5 10 3 3 2 3 $\times$ $\times$ $\times$
2.4 Theoretical Analysis
Theorem 1.
$\begin{split}\text{ERR}_{\text{out}}^{t}\leq&\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}+O(\sqrt{\frac{d}{\pi^{2}n}})+O(\sqrt{\frac{\mathrm{max}\{d,\Delta_{\xi}^{\eta^{2}}/\pi^{2}\}}{\pi^{2}(1-\pi)m}})\end{split}$ (6)
Practical implications of Therorem 1. The above theorem states that under mild assumptions, the error $ERR_{\mathrm{out}}$ is upper bounded. If the following two regulatory conditions hold: 1) the sizes of the labeled ID $n$ and wild data $m$ are sufficiently large; 2) the optimal ID risk $R_{in}^{*}$ is small, then the upper bound is mainly depended on $T$ and $t$. We further study the main error of $T$ and $t$ which we defined as $\delta(T,t)$.
Theorem 2.
$\begin{split}\delta(T,t)&=\frac{\mathrm{max}\{0,1-\Delta_{\xi}^{\eta}/\pi\}}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\leq\frac{\epsilon}{1-T/(\sqrt{2}/(2t\alpha-1))^{2}}\end{split}$ (7)
Practical implications of Therorem 2. Theorem 2 states that when the learning rate $\alpha$ is sufficiently small, the primary error $\delta(T,t)$ can approach zero if the difference $\zeta$ between the two data distributions $\mathbb{P}_{\text{wild}}$ and $\mathbb{P}_{\text{in}}$ is greater than a certain small value. Meanwhile, Theorem 2 also shows that the primary error $\delta(T,t)$ is inversely proportional to the learning rate $\alpha$ and the number of epochs ($t$). As the $t$ increases, the primary error $\delta(T,t)$ also increases, while a smaller learning rate $\alpha$ leads to a minor primary error $\delta(T,t)$. However, during training, there exists $t$ at which the error reaches its minimum.
3 Experiment
In this section, we present the main experimental results, while in Appendix F.2, we investigate feature resonance across datasets. Appendix F.3 compares scoring methods, while F.4 evaluates time efficiency. Appendix F.5 tests RSL with different GNN encoders, and F.6 examines graph-level OOD detection. Appendices F.8 and F.9 provide score distribution and node representation visualizations.
3.1 Experimental Setup
Datasets.
We conduct extensive experiments to evaluate RSL on a total of nine real-world OOD node detection datasets: six multi-category datasets, Squirrel (Rozemberczki et al., 2021), WikiCS (Mernyei and Cangea, 2020), Cora, Citeseer, Pubmed (Kipf and Welling, 2016a), and Chameleon (Rozemberczki et al., 2021) and three binary classification fraud detection datasets: YelpChi (Rayana and Akoglu, 2015), Amazon (McAuley and Leskovec, 2013), and Reddit (Kumar et al., 2019). The statistics of these datasets are summarized in Table 1. Additionally, we validate our method on four graph-level OOD detection datasets, including ENZYMES, PROTEINS (Morris et al., 2020), ClinTox, and LIPO (Wu et al., 2018). We provide detailed dataset description in the Appendix E.3.
Table 2: Unsupervised OOD detection on real-world datasets. “OOM” indicates out-of-memory, “TLE” means time limit exceeded, and “-” denotes inapplicability. Detectors with $\clubsuit$ use only node attributes, while $\spadesuit$ share RSL’s GNN backbone ( GCN ). Entropy-based methods with $\lozenge$ use true multi-category labels, and $\blacklozenge$ rely on K-means pseudo labels. Top results: 1st, 2nd.
@@ 3.1 Experimental Setup
$\text{NodeSafe}^{\spadesuit\lozenge}$ 57.82 33.57 93.64 74.81 67.93 74.85 - - - - - - - - -
$\text{GRASP}^{\spadesuit\lozenge}$ 61.38 36.95 90.77 78.46 71.52 71.08 - - - - - - - - -
@@
$\text{NodeSafe}^{\spadesuit\blacklozenge}$ 50.91 27.48 96.18 65.77 52.02 91.03 56.61 28.01 89.95 69.92 26.44 82.72 50.74 6.03 94.26
$\text{GRASP}^{\spadesuit\blacklozenge}$ 52.63 28.12 94.87 66.94 53.33 89.12 58.05 28.67 88.11 70.31 27.81 81.29 51.82 6.91 92.04
@@
Baselines.
We assess the performance of RSL against a total of twenty-one baseline methods spanning five categories: 1) Traditional outlier detection methods, including local outlier factor (Breunig et al., 2000) like LOF-KNN and MLPAE. 2) Graph-based outlier detection models, including GCN autoencoder (Kipf and Welling, 2016b), GAAN (Chen et al., 2020), DOMINANT (Ding et al., 2019), ANOMALOUS (Peng et al., 2018), and SL-GAD (Zheng et al., 2021). 3) Transformation-based outlier detection approaches, such as GOAD (Bergman and Hoshen, 2020) and NeuTral AD (Qiu et al., 2021). 4) Entropy-based detection techniques, including MSP, ODIN, OE, Energy, GKDE (Zhao et al., 2020), OODGAT (Song and Wang, 2022), GNNSafe (Wu et al., 2023), NodeSafe (Yang et al., 2025), and GRASP (Ma et al., 2024). 5) Category-free detection methods, including EnergyDef (Gong and Sun, 2024) and SSD (Sehwag et al., 2021).
Additionally, we also compare our method with graph-level approaches, including 1) graph kernel combined with a detector (Vishwanathan et al., 2010; Shervashidze et al., 2011; Neumann et al., 2016; Breunig et al., 2000; Manevitz and Yousef, 2001; Liu et al., 2008), 2) graph contrastive learning with a detector (You et al., 2020; Liu et al., 2008; Sehwag et al., 2021; Zhou et al., 2021; Liu et al., 2023), and 3) end-to-end methods (Zhao and Akoglu, 2023; Ma et al., 2022).
Details of baselines and implementation are in Appendix E.4 and E.5, respectively.
Metrics.
Following prior research on OOD node detection, we evaluate the detection performance using three widely recognized, threshold-independent metrics: AUROC ($\uparrow$), AUPR ($\uparrow$) and FPR95($\downarrow$). We provide a detailed metric description in the Appendix E.2.
Table 3: Performance comparison across methods on Cora, Citeseer, Pubmed, and Chameleon.
Method Dataset Cora Citeseer Pubmed Chameleon
[…]
MSP 70.86 84.56 67.81 82.39 87.37 68.80 85.70 57.96
Energy 67.54 85.47 88.53 72.38 93.86 54.09 88.06 59.20
KNN 90.20 70.94 83.10 72.91 89.79 64.14 93.38 57.90
ODIN 68.41 84.98 67.91 82.42 87.49 68.80 85.31 57.94
Mahalanobis 69.68 85.48 99.12 54.62 96.81 56.85 95.55 53.19
GKDE 63.71 86.27 80.42 79.94 65.48 69.92 92.93 50.14
GPN 58.45 82.93 65.68 88.13 88.61 64.13 82.25 68.20
OODGAT 94.59 53.63 62.39 84.33 88.27 58.28 94.43 59.67
GNNSafe 54.71 87.52 60.15 84.85 62.47 83.70 100.00 50.42
NodeSafe 50.32 89.11 55.71 86.16 58.07 85.11 98.76 52.19
GRASP 29.70 93.50 35.23 89.75 37.41 88.43 66.88 76.93
RSL 28.76 94.14 33.67 90.44 35.15 89.10 45.81 78.04
3.2 Main Results
Tables 2 and 3 present the main experimental results of various methods across nine public datasets. Specifically, when multi-class labels are unavailable, RSL significantly outperforms existing methods. For methods that require multi-class labels, we follow EnergyDef (Gong and Sun, 2024) by assigning pseudo-labels using K-means. On the YelpChi, Amazon, and Reddit datasets, RSL achieves average improvements of 3.01%, 7.09%, and 8.95% over the SOTA methods in terms of AUROC, AUPR, and FPR95, respectively.
When multi-class label information is available, RSL shows even more significant performance gains on heterophilic graphs. On the Squirrel, WikiCS, and Chameleon datasets, RSL achieves an average improvement of 14.93% in FPR95 over SOTA methods. This is because RSL does not rely on the homophily assumption of the graph, and thus performs well on heterophilic graphs. On homophilic graphs with multi-class labels—namely Cora, Citeseer, and Pubmed—RSL achieves performance comparable to SOTA methods. Notably, RSL does not leverage multi-class labels for training in any of the above experiments. This highlights its label-agnostic and task-independent nature, contributing to its broader applicability.
Table 4: Performance of RSL at achievable label proportion $R$ in the WikiCS dataset.
$R=0.0$ $R=0.1$ $R=0.2$ $R=0.4$ $R=0.6$ $R=0.8$ $R=1.0$
AUROC(↑) 84.01 86.41 87.46 87.89 88.35 88.57 89.07
AUPR(↑) 81.14 83.35 84.80 85.24 85.42 86.16 86.85
FPR@95(↓) 49.23 44.95 42.69 41.25 39.50 39.68 38.78
Table 5: Performance of different models on the Citeseer under varying homophily ratio $R$ values.
Model $R=0.0$ $R=0.1$ $R=0.2$ $R=0.3$ $R=0.4$ $R=0.5$
[…]
GNNSafe 60.15 84.85 52.92 86.56 47.68 87.40 45.64 89.30 42.22 90.42 37.19 91.78
NodeSafe 55.71 86.16 49.47 87.45 42.23 90.47 38.62 92.31 34.79 93.68 30.57 94.93
GRASP 35.23 89.75 30.63 91.26 27.35 94.00 16.01 96.36 13.69 97.33 10.08 98.25
RSL 33.67 90.44 28.91 91.75 16.62 94.45 9.82 96.09 7.71 97.30 4.53 98.23
How effective is resonance-based filter score $\tau$? The experimental results in the row labeled “RSL w/o classifier" of Table 2 show that using the raw resonance-based score $\tau$ to filter OOD nodes is already more effective than the SOTA method on most datasets. On the FPR95 metric, the score $\tau$ achieves an average reduction of 9.70% compared to current SOTA methods in Table 2.
How effective are the synthetic OOD nodes combined with the feature resonance score? The experimental results in the row labeled “RSL w/o $\mathcal{V}_{\mathrm{syn}}$" of Table 2 show that after removing the synthetic OOD nodes, the performance of the trained OOD classifier declined to varying degrees. This indicates that synthetic OOD nodes enhance the generalization ability of the OOD classifier, allowing it to detect more OOD nodes more accurately. It is worth noting that our synthetic OOD nodes, generated by leveraging real OOD nodes selected using $\tau$, better align with real-world OOD scenarios and, therefore, outperform EnergyDef.
Method Dataset Target Squirrel WikiCS
[…]
EnergyDef - 64.15 37.40 91.77 70.22 60.10 83.17
RSL w/o classifier A random vector 61.52 38.96 90.18 79.15 78.65 70.38
Can label information bring gains to RSL? Although RSL can perform well in scenarios without multi-class labels, we want to investigate whether multi-class labels can bring similar benefits to RSL as they do for other methods. On the WikiCS dataset, we first pre-train the representations of the training set’s ID nodes using supervised contrastive learning loss (Khosla et al., 2020) with training set labels at different proportions $R$, and then apply RSL. When $R=1.0$, it indicates that RSL, like other methods that strictly require labels, uses all the training set labels. The results in Table 4 show that as the available label proportion increases, the ID node representations of the training set are better initialized, and RSL performs better. We believe this is because when the ID node representations are well-initialized, feature resonance is more easily induced and is more pronounced.
Can graph homophily bring gains to RSL? Most existing OOD node detection methods benefit from graph homophily, so we aim to explore whether RSL can also gain from it. We conduct experiments on the Citeseer dataset under varying levels of homophily, by removing a proportion $R$ of heterophilous edges and adding the same proportion $R$ of homophilous edges. The results in Table 5 show that as graph homophily improves, the performance of RSL also improves. We believe this is because enhanced graph homophily leads to more consistent representations among ID nodes and more pronounced differences between ID and OOD node representations, thereby making feature resonance easier to induce and more strongly expressed.
How does feature resonance occur due to different target vectors? We explore micro-level feature resonance using different target vectors through experiments on Squirrel and WikiCS datasets with true $N$-category labels. Based on neural collapse theory (Papyan et al., 2020; Zhou et al., 2022), we set $N$ target vectors that form a simplex equiangular tight frame  [^3] The definition of the simplex equiangular tight frame is introduced in Appendix 6., maximizing separation. As shown in Table 6, the "True multi-label" row demonstrates the effectiveness of this approach. Interestingly, even when random labels are assigned (the "Multiple random vectors" row) or when all ID representations align with a fixed vector (the "A random vector" row), unknown ID nodes still show larger $\tau$ than unknown OOD nodes, as seen in Table 6. These results suggest that feature resonance is label-independent, stemming from intrinsic relationships between ID node representations.
@@ 3.2 Main Results
How do other OOD node selection methods perform? We aim to evaluate the performance of RSL when integrated with methods other than the resonance-based score for selecting reliable OOD nodes. To ensure fairness, we used the same parameters and selected the same number of OOD nodes. From a metric learning perspective, we computed the cosine similarity, Euclidean distance, and Mahalanobis distance between unknown nodes and the prototypes of known ID nodes, with smaller values indicating a higher likelihood of being OOD nodes. We also applied EnergyDef for OOD node selection. The results, presented in Table 7, show that, under the same conditions, the OOD nodes selected using $\tau$ are more reliable than those selected by the other methods.
4 Related Work